Your school band sold stationery for $12 a box. The profit, P, in dollars is given by P = 12x - 3640, where x is the number of boxes sold. If the band made a profit of $2600, how many boxes of stationery were sold? You Must Show All Work
Answer by ikleyn(52934)

.
Your school band sold stationery for $12 a box.
The profit, P, in dollars is given by P = 12x - 3640, where x is the number of boxes sold.
If the band made a profit of $2600, how many boxes of stationery were sold?
You Must Show All Work
~~~~~~~~~~~~~~~~~~

Write equation as you read the problem

    12x - 3640 = 2600.


Simplify and find x

    12x = 2600 + 3640 

    12x =    6240

      x =    6240/12 = 520.


ANSWER.  520 boxes of stationery were sold.

Solved, and all work is shown, to your satisfaction.